Find the surface area of a right circular cylinder that is 6cm in diameter and is 12cm in height.
Answer by prabhjyot(164)

Radius: r
Height: h
Surface area: A = 2(pi)(r^2) + 2(pi)rh
r = (6 cm) / 2 = 3 cm
h = 12 cm
A = 2(pi)[(3 cm)^2] + 2(pi)(3 cm)(12 cm) = [18(pi) cm^2] + [72(pi) cm^2] = 90(pi) cm^2